is xy-5y=1 a function such that there is one value for y for every x intercept? If so, why?
Answer by stanbon(75887)

is xy-5y=1 a function such that there is one value for y for every x intercept? If so, why?
----------------------
Solve for "y".
y(x-5) = 1
y = 1/(x-5)
-----------------------
There are no x-intercepts.
To find the x-intercept you would let y = 0 and you would get:
0 = 1/(x-5)
Then cross-multiplying you would get 0 = 1.
Obviously there are no x- intercepts.
Cheers,
Stan H.